On average, Nancy has noticed that 17 trucks pass by her apartment daily (24 hours). In order to find the probability that more than 3 trucks will pass her apartment in a 3-hour time period using the Poisson distribution, find the average number of trucks per 3 hours. Round your answer to three decimal places, if necessary.

Provide your answer below:
Answer #1

Here, 17 trucks pass in 24 hours

we need to find for 3 hours

So, average = 17 * 3/24 = 2.125


Here, λ = 2.125 and x = 3
As per Poisson's distribution formula P(X = x) = λ^x * e^(-λ)/x!

We need to calculate P(X > 3) = 1 - P(X <= 3).
P(X > 3) = 1 - (2.125^0 * e^-2.125/0!) + (2.125^1 * e^-2.125/1!) + (2.125^2 * e^-2.125/2!) + (2.125^3 * e^-2.125/3!)
P(X > 3) = 1 - (0.1194 + 0.2538 + 0.2697 + 0.191)
P(X > 3) = 1 - 0.8339 = 0.1661